# OORD: The Oxford Offroad Radar Dataset

### v1

@@ Abstract
There is a growing academic interest as well as commercial exploitation of millimetre-wave scanning radar for autonomous vehicle localisation and scene understanding. Although several datasets to support this research area have been released, they are primarily focused on urban or semi-urban environments. Nevertheless, rugged offroad deployments are important application areas which also present unique challenges and opportunities for this sensor technology. Therefore, the Oxford Offroad Radar Dataset (OORD) presents data collected in the rugged Scottish highlands in extreme weather. The radar data we offer to the community are accompanied by GPS/INS reference – to further stimulate research in radar place recognition. In total we release over of radar scans as well as GPS and IMU readings by driving a diverse set of four routes over forays, totalling approximately of rugged driving. This is an area increasingly explored in literature, and we therefore present and release examples of recent open-sourced radar place recognition systems and their performance on our dataset. This includes a learned neural network, the weights of which we also release. The data and tools are made freely available to the community at oxford-robotics-institute.github.io/oord-dataset
@@ I Introduction
Compared to cameras and LiDARs, radar systems offer distinct advantages, including extended-range perception and resilience to lightning and various weather conditions. Full reviews of millimetre-wave radar applications in autonomous vehicle tasks are available in e.g. [1, 2]. In brief, scanning radar-based motion estimation [3, 4, 5, 6, 7, 8, 9, 10], pose estimation [11, 12, 13], place recognition [14, 15, 11, 16, 17, 18, 19, 20, 21, 22, 23], and full simultaneous localisation and mapping (SLAM) [24, 25] are increasingly popular in the literature. This dataset is designed to support that work, and to prompt an extension into naturalistic, rugged environments, which have not yet been explored.
@@ II-A Radar datasets
Datasets featuring a similar class of millimetre-wave scanning radar include Marulan [26] by Peynot et al, which is furthermore also collected in natural environments, as is ours. This data, however, only has range of up to . Several recent radar datasets use the same radar class – from the same manufacturer111Navtech Radar: navtechradar.com – as used in our work, including the Oxford Radar RobotCar Dataset [27] by Barnes et al, MulRan [14] by Kim et al, RADIATE [28] by Sheeny et al, Boreas [29] by Burnett et al, and OSDaR23 [30] by Tagiew et al. The Oxford Radar RobotCar Dataset [27] is purely urban and features many repeat traversals, useful in investigating place recognition. The ground truth provided in [27], however, is prepared for radar odometry rather than place recognition. MulRan [14] presents more diversity in scenery, having been collected in several distinct parts of Daejeon, South Korea. We similarly attend to generalisation-sensitive training/testing data and collect at many sites, albeit offroad and more widespread across a country rather than a city. Our repeat traversals are also more numerous than for MulRan. RADIATE [28], presenting vehicle detection and tracking ground truth, is similarly lacking in overlapping traversals and not geared towards place recognition, although a diversity in weather makes this dataset interesting for exploring this sensor’s robustness. Closest to our work is Boreas by Burnett et al [29], who collect multi-season data including a Navtech CIR304-H scanning radar. Our work is in part inspired by this, with ours featuring highly non-planar offroad driving as opposed to purely urban.
@@ II-B Radar place recognition
This area is therefore of increasing interest, but far from mature, and this dataset should therefore contribute to establishing useful benchmarks for the community.
@@ III Off-road routes & collection sites
| Route | Distance | # Radar | # GPS | Size |
| Bellmouth | ||||
| Hydro | ||||
| Maree | ||||
@@
Figs. 2 and 6 show a detailed overview of the datasets we provide, at various areas of the Ardverikie Estate. Data has been gathered in the region encompassing Loch Laggan and Lochan na h-Earba. The position of these lochs can be particularly well understood by referring to the caption of Fig. 5(b).
Each of Figs. 5(b), 1(a), 5(a) and 1(b) is arranged as follows: Top Left: Photograph of each dataset collection location (handheld camera, not on-board imagery). Top Middle: GPS trace of route driven. Top Right: Ground truth match matrix for a pair of trajectories at that site. Bottom: Example radar scans from various points along the route (Cartesian).
@@ III-A Route 1: Bellmouth
This is a route. The route is completely landlocked, heading away from and then back towards Loch Laggan, see Fig. 1(a) (Top Middle. The ascent and descent are reflected by a somewhat symmetrical elevation change in Fig. 3 – climbing for the first half of the route, and descending for the second half. The vehicle is always on a gravel track. Figs. 3(b), 3(a), 3(d) and 3(c) show images from the vehicle’s perspective from four outings along this route. The first two these outings is prior to any snow falling in the Ardverikie Estate area (Figs. 3(a) and 3(b)). The next two of these outings show significant snow dump (Figs. 3(c) and 3(d)). The outings are in the same direction, as per the ground truth matrix in Fig. 5(b) (Top Right), but within each outing itself we double back (c.f. the off diagonal matches in three parts of this matrix).
@@ III-B Route 2: Hydro
This is a route. As shown in Fig. 3 this route reaches the highest elevation of all or our collection sites – ascending a gravel track to maximum elevation and then descending along the same route. For parts of the route, we travel along a mountain river, as per Fig. 1(b) (Top Middle). The snow coverage is significant (Figs. 3(e), 3(f) and 3(g)), but not as thick as for Maree below.
@@ III-C Route 3: Maree
This is a route. The route is mountainous, in the elevated area to the south of Loch Laggan, see Fig. 5(a) (Top Middle). There are two ascents and descents as can be seen in Fig. 3 – with a descent halfway through the journey, and another at the end of the journey. The vehicle is primarily on a gravel track. Figs. 3(h) and 3(i) show images from the vehicle’s perspective from two outings along this route. Both of these feature heavy snow, with a completely covered track. In Fig. 3(i) the vehicle is driving in total darkness (with no artificial illumination beyond the vehicle’s spotlights as this is a very remote and isolated natural collection site). The outings are in the same direction, as per the ground truth matrix in Fig. 5(a) (Top Right), but within each outing itself we double back on parts of the route and return to the starting point along the same route we took to begin the route (c.f. the off diagonal matches in three parts of this matrix).
@@ III-D Route 4: Two Lochs
This is a route across the entire estate. The route is alongside two lochs (Loch Laggan and Lochan na h-Earba), see Fig. 5(b) (Top Middle, where the lochs are on top of and the bottom of the image, respectively). Referring to the route in an anti-clockwise sense, while going alongside Loch Laggan the vehicle is on tarmac – it is otherwise on unpaved road for the rest of the route. Figs. 3(k) and 3(j) show images from the vehicle’s perspective of our two outings – these are completed after snow has fallen (in previous days), so melted somewhat, but while it is still very visible on the ground etc. The outings are in the opposite direction as shown by the off-diagonal matches in the ground truth matrix in Fig. 5(b) (Top Right). Note that the red elevation profile in Fig. 3 corresponds to the foray shown in Fig. 5(b) (Top Middle), and is very flat at approximately as the vehicle traverses alongside Lochan na h-Earba.
@@ III-E Summary: Datasets provided
(Dataset 11)[DA2]: 2021-11-29-12-19-16 (Two Lochs)
These are shown in Fig. 4. We note that, beyond the weather and illumination effects discussed for each route above, some forays have lens adherents on the on-board camera (e.g. Figs. 3(c) and 3(d) in particular), motivating the use of radar in this sort of domain and for the localisation task.
@@ IV Sensors & dataset format
At oxford-robotics-instiute.github.io/oord-dataset we provide for download time-stamped synchronised radar and GPS/INS.
@@ IV-A Radar
IV-B GPS/INS
@@ V-A Downloads & data loaders
For working with driven sequences, we provide a basic iterable Dataset class which returns each radar scan paired with its closest (in a timestamp sense) GPS/INS reading (Sec. IV-B).
@@
V-B Radar place recognition evaluation
We evaluate place recognition performance by comparing “distance matrices” in the ground truth space given by GPS, as shown by example in Figs. 2 and 6 (Top Right) to distance matrices composed of e.g. pairwise distances between vectors representing radar scans (see Sec. VI) or some other similarity measurement. These matrices have heights given by the number of radar scans in a “query” foray and widths similarly for a “reference” foray. Highlighted regions are such that those radar scan pairs have locations given by GPS which are within (this threshold being configurable in our SDK). In these, interesting features to note are the off-diagonal revisits, which represent traversals in the opposite direction. This is especially important in the radar modality as the sensor has a field-of-view.
Then, Recall@N refers to the percentage of query scans which result in at least one successful localisation (i.e. resulting in a returned location within of the query location) when accepting as candidates all of the nearest neighbours in the embedding space. This is a commonly used metric and threshold, as in visual place recognition [36, 37]. Note that in Figs. 2 and 6 the true positives are “thickened” for visualisation purposes, to , but we experiment in Sec. VII with .
For this we use a “difference matrix” in Fig. 5 with the embedding distances (Euclidean) between live and map features. Localisation is successful if the nearest neighbour for a query embedding (rows) is a reference embedding (column) which is in truth (c.f. the ground truth matrices in Fig. 1(a) Top Right) is close to the query location in physical space.
@@ V-C Radar place recognition with neural networks
As will be discussed below, we evaluate a family of learned methods on our dataset. This includes a pretrained and frozen backbone feature extractor along with a pooling layer which is fine-tuned on radar data from this dataset, and we release these weights at huggingface.co/mttgdd/oord-models , with this being, to our knowledge, the first open-sourced learned radar place recognition system.
@@ VI Examples
We demonstrate the utility of our dataset as a novel radar place recognition benchmark using the evaluation method described in Sec. V-B. We also show that the GPS/INS described in Sec. IV-B can be put to good use in supervising learned radar place recognition methods.
Therefore, we demonstrate implementations of the following methods with our dataset:
@@
Sec. VII presents a comprehensive comparison of these example baselines over all pairs of trajectories from our dataset.

### v2

@@ Abstract
There is a growing academic interest as well as commercial exploitation of millimetre-wave scanning radar for autonomous vehicle localisation and scene understanding. Although several datasets to support this research area have been released, they are primarily focused on urban or semi-urban environments [1, 2, 3, 4]. Nevertheless, rugged offroad deployments are important application areas which also present unique challenges and opportunities for this sensor technology. Therefore, the Oxford Offroad Radar Dataset (OORD) presents data collected in the rugged Scottish highlands in extreme weather. The radar data we offer to the community are accompanied by GPS/INS reference – to further stimulate research in radar place recognition. In total we release over of radar scans as well as GPS and IMU readings by driving a diverse set of four routes over forays, totalling approximately of rugged driving. This is an area increasingly explored in literature, and we therefore present and release examples of recent open-sourced radar place recognition systems and their performance on our dataset. This includes a learned neural network, the weights of which we also release. The data and tools are made freely available to the community at oxford-robotics-institute.github.io/oord-dataset
@@ I Introduction
Compared to cameras and LiDARs, radar systems offer distinct advantages, including extended-range perception and resilience to lightning and various weather conditions. In more detail, scanning radar of the type we use in this work does not give 3D information, and is suspect to complex measurement artefacts (speckle noise, ghost reflections, etc), but is extremely robust to atmospheric effects and low-light. Cameras are information-rich and highly discriminative, but are highly susceptible to appearance variation, e.g. low-light. LiDARs give detailed scans, often in 3D, and emit their own light so are suitable for low-light operation, but are often limited in range and are nevertheless susceptible to particulate effects, e.g. snow or dust. Full reviews of millimetre-wave radar applications in autonomous vehicle tasks are available in e.g. [5, 6]. In brief, scanning radar-based motion estimation [7, 8, 9, 10, 11, 12, 13, 14], pose estimation [15, 16, 17], place recognition [18, 19, 15, 20, 21, 22, 23, 24, 25, 26, 27], and full simultaneous localisation and mapping (SLAM) [28, 29] are increasingly popular in the literature. This dataset is designed to support that work, and to prompt an extension into naturalistic, rugged environments, which have not yet been explored.
@@
Drone image of route.
GPS trace (red).
Ground truth matrix.
Example scan #1.
Example scan #2.
Example scan #3.
Example scan #4.
DSLR image from route.
GPS trace (red).
Ground truth matrix.
Example scan #1.
Example scan #2.
Example scan #3.
Example scan #4.
DSLR image from route.
GPS trace (red).
Ground truth matrix.
Example scan #1.
Example scan #2.
Example scan #3.
Example scan #4.
DSLR image from route.
GPS trace (red).
Ground truth matrix.
Example scan #1.
Example scan #2.
Example scan #3.
Example scan #4.
@@ II-A Radar datasets
Datasets featuring a similar class of millimetre-wave scanning radar include Marulan [30] by Peynot et al, which is furthermore also collected in natural environments, as is ours. This data, however, only has range of up to . Several recent radar datasets use the same radar class – from the same manufacturer111Navtech Radar: navtechradar.com – as used in our work, including the Oxford Radar RobotCar Dataset [31] by Barnes et al, MulRan [18] by Kim et al, RADIATE [32] by Sheeny et al, Boreas [33] by Burnett et al, and OSDaR23 [34] by Tagiew et al. The Oxford Radar RobotCar Dataset [31] is purely urban and features many repeat traversals, useful in investigating place recognition. The ground truth provided in [31], however, is prepared for radar odometry rather than place recognition. MulRan [18] presents more diversity in scenery, having been collected in several distinct parts of Daejeon, South Korea. We similarly attend to generalisation-sensitive training/testing data and collect at many sites, albeit offroad and more widespread across a large area of countryside/wilderness rather than a city. Our repeat traversals are also more numerous than for MulRan. RADIATE [32], presenting vehicle detection and tracking ground truth, is similarly lacking in overlapping traversals and not geared towards place recognition, although a diversity in weather makes this dataset interesting for exploring this sensor’s robustness. Closest to our work is Boreas by Burnett et al [33], who collect multi-season data including a Navtech CIR304-H scanning radar. Our work is in part inspired by this, with ours featuring highly non-planar offroad driving as opposed to purely urban.
@@ II-B Radar place recognition
This place recognition capability can be crucial for simultaneous localisation and mapping (SLAM), where reinitialisation is essential for correcting errors and maintaining accuracy over time. SLAM is a newly developed area for scanning radar in autonomous vehicle operations, as per [28, 29, 37], and so robust reinitialisation through radar place recognition is an important research topic. This area is therefore of increasing interest, but far from mature, and this dataset should therefore contribute to establishing useful benchmarks for the community.
@@ III Off-road routes & collection sites
Figs. 3 and 5 show a detailed overview of the datasets we provide, at various areas of the Ardverikie Estate. Data has been gathered in the region encompassing Loch Laggan and Lochan na h-Earba. The position of these lochs can be particularly well understood by referring to the caption of Fig. 5(b).
Each of Figs. 5(b), 3(a), 5(a) and 3(b) is arranged as follows: Top Left: Photograph of each dataset collection location (handheld camera, not on-board imagery). Top Middle: GPS trace of route driven. Top Right: Ground truth match matrix for a pair of trajectories at that site. Bottom: Example radar scans from various points along the route (Cartesian).
| Route | Distance | # Radar | # Camera | # GPS | Size |
| Bellmouth | |||||
| Hydro | |||||
| Maree | |||||
@@ III-A Route 1: Bellmouth
This is a route. The route is completely landlocked, heading away from and then back towards Loch Laggan, see Fig. 3(a) (Top Middle. The ascent and descent are reflected by a somewhat symmetrical elevation change in Fig. 6 – climbing for the first half of the route, and descending for the second half. The vehicle is always on a gravel track. Figs. 7(b), 7(a), 7(d) and 7(c) show images from the vehicle’s perspective from four outings along this route. The first two these outings is prior to any snow falling in the Ardverikie Estate area (Figs. 7(a) and 7(b)). The next two of these outings show significant snow dump (Figs. 7(c) and 7(d)). The outings are in the same direction, as per the ground truth matrix in Fig. 5(b) (Top Right), but within each outing itself we double back (c.f. the off diagonal matches in three parts of this matrix).
@@ III-B Route 2: Hydro
This is a route. As shown in Fig. 6 this route reaches the highest elevation of all or our collection sites – ascending a gravel track to maximum elevation and then descending along the same route. For parts of the route, we travel along a mountain river, as per Fig. 3(b) (Top Middle). The snow coverage is significant (Figs. 7(e), 7(f) and 7(g)), but not as thick as for Maree below.
@@ III-C Route 3: Maree
This is a route. The route is mountainous, in the elevated area to the south of Loch Laggan, see Fig. 5(a) (Top Middle). There are two ascents and descents as can be seen in Fig. 6 – with a descent halfway through the journey, and another at the end of the journey. The vehicle is primarily on a gravel track. Figs. 7(h) and 7(i) show images from the vehicle’s perspective from two outings along this route. Both of these feature heavy snow, with a completely covered track. In Fig. 7(i) the vehicle is driving in total darkness (with no artificial illumination beyond the vehicle’s spotlights as this is a very remote and isolated natural collection site). The outings are in the same direction, as per the ground truth matrix in Fig. 5(a) (Top Right), but within each outing itself we double back on parts of the route and return to the starting point along the same route we took to begin the route (c.f. the off diagonal matches in three parts of this matrix).
@@ III-D Route 4: Two Lochs
This is a route across the entire estate. The route is alongside two lochs (Loch Laggan and Lochan na h-Earba), see Fig. 5(b) (Top Middle, where the lochs are on top of and the bottom of the image, respectively). Referring to the route in an anti-clockwise sense, while going alongside Loch Laggan the vehicle is on tarmac – it is otherwise on unpaved road for the rest of the route. Figs. 7(k) and 7(j) show images from the vehicle’s perspective of our two outings – these are completed after snow has fallen (in previous days), so melted somewhat, but while it is still very visible on the ground etc. The outings are in the opposite direction as shown by the off-diagonal matches in the ground truth matrix in Fig. 5(b) (Top Right). Note that the red elevation profile in Fig. 6 corresponds to the foray shown in Fig. 5(b) (Top Middle), and is very flat at approximately as the vehicle traverses alongside Lochan na h-Earba.
@@ III-E Summary: Datasets provided
(Two Lochs #2)[DA2]: 2021-11-29-12-19-16
These are shown in Fig. 7. We note that, beyond the weather and illumination effects discussed for each route above, some forays have lens adherents on the on-board camera (e.g. Figs. 7(c) and 7(d) in particular), motivating the use of radar in this sort of domain and for the localisation task.
@@ IV Sensors & dataset format
At oxford-robotics-institute.github.io/oord-dataset we provide download time-stamped synchronised radar and GPS/INS.
@@ IV-A Radar
IV-B Camera
Our vehicle was also equipped with a FLIR Blackfly USB3 Vision camera222flir.com/products/blackfly-s-usb3 which captures images at resolution over a sensing area and with of dynamic range.
Some example images are shown in Fig. 7, and we present experimental results comparing vision to radar in Sec. VII. Note that in Tab. II the camera framerate has been downsampled to more approximately match the radar frame rate (and mitigate download sizes).
This will be useful for sensor fusion research such as [40], where our radar and camera are timestamp-synchronised.
IV-C GPS/INS
@@
IV-D Calibration
Extrinsic calibration between the sensors is done by reference to a detailed CAD model of the vehicle and sensor suite. These are included in the SDK (Sec. V) as an extrinsics.yaml file specifying the relative six degree-of-freedom pose between each of the GPS/INS (Sec. IV-C)
which is to say that the GPS unit is approximately behind the radar with axes aligned to the radar. Then, for the camera (Sec. IV-B) to radar
which is to say that the camera is further behind and to the right of the radar with axes offset by two rotations. Finally, the radar (Sec. IV-A, as the main sensor focused on in this work) is at , i.e. is used as the datum/origin of the frame of reference.
@@ V-A Downloads & data loaders
For working with driven sequences, we provide a basic iterable Dataset class which returns each radar scan paired with its closest (in a timestamp sense) GPS/INS reading (Sec. IV-C).
@@
V-B Radar place recognition & evaluation
In place recognition it is common to represent the input image or scan by some descriptor vector. With map descriptors at places indexed by and of dimension and a query descriptor , place recognition can be stated as
where we decide that we are in place given that the descriptor for that place is closest to our query descriptor (i.e. “place recognition”). Alternatively, instead of Eq. 3, we may search for a set of nearest neighbours.
We evaluate place recognition performance by comparing “distance matrices” in the ground truth space given by GPS, as shown by example in Figs. 3 and 5 (Top Right) to distance matrices composed of e.g. pairwise distances between vectors representing radar scans (see Sec. VI) or some other similarity measurement. These matrices have heights given by the number of radar scans in a “query” foray and widths similarly for a “reference” foray.
Specifically, with respect to Eq. 3 above we have black and white distance matrices as inin Figs. 3 and 5 (Top Right) given by
where is the UTM northing/easting GPS coordinates corresponding to the query/localisation trajectory scan and reference/map trajectory scan , with and with and as the number of frames in these trajectories, respectively. Similarly, we have
with and embeddings of the query/reference radar scans at those same locations . After reducing to its binarised form with in each row indicating the nearest neighbour (or, generally, a set of s indicate some nearest neighbours) as per Eq. 3, we expect for good place recognition that (with a numerical accuracy defined as Recall@N below).
Highlighted regions are such that those radar scan pairs have locations given by GPS which are within (this threshold being configurable in our SDK). In these, interesting features to note are the off-diagonal revisits, which represent traversals in the opposite direction. This is especially important in the radar modality as the sensor has a field-of-view.
Then, Recall@N refers to the percentage of query scans which result in at least one successful localisation (i.e. resulting in a returned location within of the query location) when accepting as candidates all of the nearest neighbours in the embedding space. This is a commonly used metric and threshold, as in visual place recognition [41, 42].
Consider the black and white matrices shown in Figs. 5(b), 3(a), 5(a) and 3(b) (Top Right). Each of these has, as per Eq. 4 and Eq. 5, a number of rows equal to the number of scans (with some examples in the bottom row of Figs. 3 and 5) collected along a query route (shown by the red GPS traces in Figs. 3 and 5) and number of columns equal to the number of scans collected along a reference route – i.e. the reference route acts as a map and the query route is localised against it. Each cell in these matrices are thus a pair of scans from a pair of locations. The black background means that the distance between those locations is and white entries indicate that the pair of locations is . Note that in Figs. 3 and 5 the true positives are “thickened” for visualisation purposes, to , but we experiment in Sec. VII with .
For this we use a “difference matrix” in Fig. 8 with the embedding distances (Euclidean) between live and map features. Not that the “grids” apparent in Figs. 8(a) and 8(b) are artefacts resulting from embeddings representing scans being very similar to many such embeddings in the map. These are evident as “blocks” of blue in the difference matrix. Within this block, all embeddings are very similar (where yellow indicates embeddings being more dissimilar). This phenomena is therefore known as “perceptual ambiguity” or “aliasing”. This indicates poor representational capability of the method, where we want only scans from the same location to be similar – see Figs. 8(c) and 8(d), in which only the diagonal and lines perpendicular to the diginal are blue, with no such gridlike structure. Localisation is successful if the nearest neighbour for a query embedding (rows) is a reference embedding (column) which is in truth (c.f. the ground truth matrices in Fig. 3(a) Top Right) is close to the query location in physical space.
@@ V-C Radar place recognition with neural networks
As introduced in Sec. V-B, radar place recognition is concerned with learning discriminative and compressed representations of radar scans, to maximise similarity between likewise places and minimise similarity otherwise. However, for this and similar sensors – despite their inherent immunity to appearance change – the beam is not narrow, power returns are perturbed by noises (e.g. speckle) and artefacts (multi-path reflections, etc). However, radar data – despite being complex in this way – contains highly informative information. While it is difficult to manually/visually inspect radar data for this information, it has been found in [19, 35, 27, 23] that neural networks are able to learn to encode very informative feature representations that are germane to the task of place recognition. As will be discussed below, we evaluate a family of learned methods on our dataset. This includes a pretrained and frozen backbone feature extractor along with a pooling layer which is fine-tuned on radar data from this dataset, and we release these weights at huggingface.co/mttgdd/oord-models, with this being, to our knowledge, the first open-sourced learned radar place recognition system.
@@ VI Examples
We demonstrate the utility of our dataset as a novel radar place recognition benchmark using the evaluation method described in Sec. V-B. We also show that the GPS/INS described in Sec. IV-C can be put to good use in supervising learned radar place recognition methods. Therefore, we demonstrate implementations of the following methods with our dataset , including four methods, LABEL:B1–LABEL:B4, which are either not fine-tuned at all (by virtue of the method itself) or are fine-tuned on this data as well as four methods which are totally pre-trained, LABEL:B5–LABEL:B9, i.e. not fine-tuned on this new dataset.
Non-learned radar-specific methods
@@
The pretrained networks are selected as they have been applied to visual place recognition in other work – e.g. MobileNet by Wu et al. [51], AlexNet by Sünderhauf et al. [52], GoogleNet by Yue et al. [53], VGG by Arandjelovic et al. [41], and ResNet by Yu et al. [54] – and it is therefore important to apply them to place recognition in the radar domain. Sec. VII presents a comprehensive comparison of these example baselines over all pairs of trajectories from our dataset. A comparison of the computational efficiency of these various baselines is outside of the scope of this work, but the interested reader is referred to [55] (Figure 2, page 2) for an in-depth comparison of inference speeds and memory requirements for the pretrained networks LABEL:B5–LABEL:B9 and to [39] for a comparison of the radar-specific methods, LABEL:B1–LABEL:B3, where Open-RadVLAD is shown to be more computationally efficient than RaPlace.